OpenStudy (anonymous):

A six-sided fair die is rolled 4 times in a row. The probability of getting a 4 only on the last trial is

OpenStudy (boldjon):

you got to work out the scenario its not 4, not 4 , not 4, 4

OpenStudy (boldjon):

P(4) = 1/6 , P(not 4) = 5/6 multiply these probabilities

OpenStudy (boldjon):

the throws are independent of each other so the probs are multiplied required probability = 5/6 * 5/6 * 5/6 * 1/6

OpenStudy (boldjon):

i'm trying to help u here hailey at least work with me